OpenStudy (anonymous):

what is 4√7x + √7x simplify Too dumb to remember what it is

OpenStudy (solomonzelman):

well you know that \(\Large\color{blue}{ \sf \sqrt{ a }+ 3\sqrt{ a } = (3+1)\sqrt{ a }=4 \sqrt{ a }}\)

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \sf 4\sqrt{7x} +\sqrt{7x}= }\)\(\Large\color{blue}{ \sf 4\sqrt{7x} +1\sqrt{7x}= }\)\(\Large\color{blue}{ \sf (4+1)\sqrt{7x} = ?}\)

OpenStudy (anonymous):

5√7x?

OpenStudy (solomonzelman):

YES ! YES ! you got this!

OpenStudy (anonymous):

lol thanks man

OpenStudy (solomonzelman):

You welcome :)